Aunt Karen is eight years younger than twice the age of Mary. Their ages add up to twenty-eight. Find the ages of Aunt Karen and Mary.

Answer by Maths68(1474)


Let
Present age of Aunt Karen = k
Present age of Mary = m

Aunt Karen is eight years younger than twice the age of Mary
k+8=2m
k=2m-8..........(1)
Their ages add up to twenty-eight.
k+m=28.........(2)
Put the value of k from (1) to (2)
2m-8+m=28
3m=28+8
3m=36
3m/3=36/3
m=12
Put the value of m in (1)
k=2m-8..........(1)
k=2(12)-8
k=24-8
k=16

Present age of Aunt Karen = k = 16
Present age of Mary = m = 12